the minimum distance required to stop a car moving at 35 mi/h is 42.6 ft what is the minimum stopping distance for the same car moving at 70 mi/h


1
Expert's answer
2022-01-10T09:09:28-0500

Given:

"v_1=35\: \rm mi/h"

"d_1=\rm 42.6 \: ft"

"v_2=70\: {\rm mi/h}=2v_1"


The energy-work theorem says

"\frac{mv^2}{2}=Fd"

Hence, the stopping distance

"d\sim v^2"

So

"d_1\sim v_1^2\\
d_2\sim v_2^2=(2v_1)^2=4d_1\\=4*42.6\rm \:ft=170.4\: ft"